OpenStudy (anonymous):

The odds are 2 to 1 that, when A play B in tennis, A wins. Suppose that A and B play two matches. What is the probability that A wins at least one match?

OpenStudy (anonymous):

The simple events are E1:{AA} E2:{AB} E3:{BA} E4:{BB}

OpenStudy (anonymous):

for example AB means A won the first mach and B wins the second match

OpenStudy (anonymous):

My books says that these events are not equally proable, since A wins 2 to 1 eveytime A and B play. They assign the follwing proablilites: P(E1):=4/9, P(E2)=2/9, P(E3)=2/9 and P(E4):=1/9

OpenStudy (anonymous):

My question is: How did they get this proablilites?

OpenStudy (anonymous):

P(A wins at least 1 match) = 1-P(A loses both matches) =1-1/9 = 8/9

OpenStudy (anonymous):

okay, but how did they come up with the probablilityes?